OpenStudy (anonymous):

Determine which polynomial is a difference of two squares. x2 + 14 x2 − 14 x2 + 49 x2 − 49

OpenStudy (anonymous):

x^2-49

OpenStudy (campbell_st):

the keys are 49 = 7^2, x^2=(x)^2 and difference is subtraction... two squares so only 2 terms